# Statistical Inference for Policy Evaluation with Temporal Difference Learning

### v3

@@ 3.4 Statistical inference
Combining the results from Sections 3.2 and 3.3, we can now propose a general methodology for constructing convex confidence sets for $\bm{\theta}^{\star}$ with guaranteed finite sample coverage. Based on the TD estimator $\overline{\bm{\theta}}_{T}$ (9) and the variance estimator $\widehat{\bm{\Lambda}}_{T}$ (15), and given any tolerance level $\delta\in(0,1)$, it is enough to pick any convex set $\mathcal{C}_{\delta}$ that has probability at least $1-\delta$ under the (random!) Gaussian distribution $\mathcal{N}(\overline{\bm{\theta}}_{T},\frac{\widehat{\bm{\Lambda}}_{T}}{T})$. That is, ssattisfying, the condition
@@
Natural choices of $\mathcal{C}_{\delta}$ are (i) hyperctangles, leading to simultaneous confidence intervals of the form
$\mathcal{C}_{\delta}=\prod_{j=1}^{d}[(\overline{\bm{\theta}}_{T})_{j}-R/\sqrt{T},(\overline{\bm{\theta}}_{T})_{j}+R/\sqrt{T}],$ (20)
where $R>0$ is the quantile of the $\ell_{\infty}$ norm of a centered Gaussian random vectors with covariance $(\widehat{\bm{\Lambda}}_{T}/{T})$, which can be calculated via simulations; and (ii) ellipsoids or balls, for example
@@
The next result, whose proof is in Appendix B.4, derives a coverage rate for $\mathcal{C}_{\delta}$, which converges to the nominal level of $1-\delta$ at a rate of order $O(T^{-1/3})$, as a function of $T$ alone.
@@
We highlight two important features of the above result. First, the coverage guarantee holds uniformly over any choice of the convex set $\mathcal{C}_{\delta}$ meeting condition (19), including the common choices like (20) and (21). Secondly, the size (i.e., volume) of the confidence sets vanishes at the rate of $O(T^{-1/2})$, keeping all the other variable constant, suggesting that, in this regard, there is little or no room for improvements, at least in terms of rates.
@@ 4 Simulations
We illustrate our theoretical results with some numerical experiments. [^3] The code for our experiments can be found at https://www.dropbox.com/scl/fo/tbw256dlx4gzxzgb4t3hq/AOUaiLFj79iSnc29dFTPpvQ?rlkey=nsgwlcbfx0b73uebwv8xlovag&st=1axbxaa1&dl=0 In order to test our proposed framework for statistical inference, we consider a member of a family of MDPs that are designed to be difficult to distinguish among each other. The specifics of the MDP are the same as described in Li et al. (2024), and we refer readers to Appendix D for the construction details. With $10,000$ independent trials, we examine the high-probability convergence of the $\ell_{2}$ norm of TD estimation errors, the rate by which they converge to their asymptotic Gaussian distributions, and the precision of our estimator for the covariance matrix. Figure 2 reveals that the empirical distribution of TD estimation error converges to the asymptotic Gaussian distribution slowly when $\alpha$ is set to ${1}/{2}$ as suggested by Samsonov et al. (2024). Notice that, for computational simplicity, the difference bewteen the empirical distribution of $\sqrt{T}\overline{\bm{\Delta}}_{T}$ and the asymptotic distribution $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$ is measured by
$d(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})):=\max_{1\leq j\leq d}\sup_{x\in\mathbb{R}}\left|\mathbb{P}(\sqrt{T}(\overline{\bm{\Delta}}_{T})_{j}\leq x)-\mathbb{P}((\bm{\Lambda}^{\star\frac{1}{2}}\bm{z})_{j}\leq x)\right|$
instead of the convex distance we invoked in the theoretical results.
Figure 2: Distance between empirical and asymptotic distributions of the TD estimation error.
@@ 5 Discussion
We develop high-probability convergence guarantees and Berry-Esseen bounds for TD learning with linear function approximation and $i.i.d.$ samples in a non-asymptotic manner. We also propose an efficient way to construct a covariance estimator and derive its consistency rates. These results lead to a practicable methodology for constructing convex confidence sets for the target parameter, with its size vanishing at a parametric rate $O(T^{-1/2})$. An important direction for future research is to investigate whether the order $O(T^{-1/3})$ for the convergence speed of coverage rate in Theorem 3.4 are tight. Another interesting extension would be to derive specialized Berry-Esseen bounds for the sub-class of hyper-rectangles, where one would expect to find a better dependence on the dimension $d$ of the feature space, possibly of poly-logarithmic, as opposed to polynomial, order.
We believe our analysis and results can be adapted or extended to cover other iterative algorithms, such as stochastic gradient descent and SA methods. Recent analyses (see, e.g., Sheshukova et al., 2025; Samsonov et al., 2024) have demonstrated Berry-Esseen bounds for the convex distance (using an estimated variance and assuming i.i.d. data), as well as valid bootstrap schemes.
Acknowledgments and Disclosure of Funding
@@ A.1 Useful algebraic facts
For any complex number $z\in\mathbb{C}$, we use $\mathsf{Re}(z)$ to denote its real part. If a matrix $\bm{M}$ is positive (semi-)definite, i.e. $\lambda_{\min}(\bm{M})>(\geq)0$, we write $\bm{M}\succ(\succeq)\bm{0}$; when two matrices $\bm{X}$ and $\bm{Y}$ satisfy $\bm{X}-\bm{Y}\succ(\succeq)\bm{0}$, we write $\bm{X}\succ(\succeq)\bm{Y}$ or equivalently $\bm{Y}\prec(\preceq)\bm{X}$. For a vector $\bm{x}\in\mathbb{R}^{d}$, we use $\{x_{i}\}_{1\leq i\leq d}$ to denote its entries; for a matrix $\bm{X}\in\mathbb{R}^{m\times n}$, we use $\{X_{ij}\}_{1\leq i\leq m,1\leq j\leq n}$ to denote its entries. The vectorization of matrix $\bm{X}\in\mathbb{R}^{m\times n}$, denoted as $\textbf{vec}(\bm{X})$, is defined as a vector in $\mathbb{R}^{mn}$ with entries $[\textbf{vec}(\bm{X})]_{(i-1)m+j}=\bm{X}_{ij}$ for every $1\leq i\leq m$ and $1\leq j\leq n$. For any matrices $\bm{X}\in\mathbb{R}^{m\times n}$, and $\bm{Y}\in\mathbb{R}^{p\times q}$, their Kronecker product, denoted as $\bm{X}\otimes\bm{Y}$, is a matrix of the size of $mp\times nq$, and can be written in blocked form as
$\bm{X}\otimes\bm{Y}=\begin{bmatrix}X_{11}\bm{Y}&X_{12}\bm{Y}&\ldots&X_{1n}\bm{Y}\\ X_{21}\bm{Y}&X_{22}\bm{Y}&\ldots&X_{2n}\bm{Y}\\ \ldots&\ldots&\ldots&\ldots\\ X_{m1}\bm{Y}&X_{m2}\bm{Y}&\ldots&X_{mn}\bm{Y}\end{bmatrix}.$
With subscripts, it can be represented as $(\bm{X}\otimes\bm{Y})_{(i-1)m+j,(k-1)p+\ell}=X_{ik}Y_{j\ell}$. For a matrix sequence $\{\bm{F}_{t}\}_{t\in\mathbb{N}}$, a number sequence $\{f_{t}\}_{t\in\mathbb{N}}$ and a fixed matrix $\bm{X}$, if there exists a constant $\widetilde{C}$ independent of $t$ and $\bm{X}$ such that
@@
The following lemma, due to Chen et al. (2020)[Lemma 4.1], bounds the difference between the inverse matrices and would be useful for our analysis of the plug-in variance estimator.
@@
00 0 $\leq\mathbb{E}[f^{2}(s)]-2\mathbb{E}[f(s)\bm{\phi}^{\top}(s)\bm{y}]+\bm{y}^{\top}\mathbb{E}[\bm{\phi}(s)\bm{\phi}(s)^{\top}]\bm{y}$
@@
<$\overset{(i)}{<}C+C\left(\frac{k+1}{k}\right)^{\nu-\alpha}\cdot\frac{\nu-\alpha}{k+1}-\beta C(k+1)^{-\alpha}\left(\frac{k+1}{k}\right)^{\nu-\alpha}+(k+1)^{-\alpha}$
<$\overset{(ii)}{<}C-\frac{\beta}{2}C(k+1)^{-\alpha}\left(\frac{k+1}{k}\right)^{\nu-\alpha}+(k+1)^{-\alpha}$
<$\overset{(iii)}{<}C-(k+1)^{-\alpha}\left(\frac{k+1}{k}\right)^{\nu-\alpha}+(k+1)^{-\alpha}<C.$
@@
where $Z_{1}=1^{\alpha}X_{1}-\frac{1}{\beta}=1-\frac{1}{\beta}$. The proof now boils down to showing that both $I_{1}$ and $I_{2}$ converges to 00 as $t\to\infty$. On one hand, $I_{1}$ is bounded by
@@
=$\overset{(i)}{=}\frac{1}{\beta}\left(1+\sum_{i=t}^{\infty}\prod_{k=t+1}^{i+1}(1-\beta k^{-\alpha})-\sum_{j=t}^{\infty}\prod_{k=t+1}^{j+1}(1-\beta k^{-\alpha})\right)=\frac{1}{\beta}.$ (45)
@@
($-\underset{I_{23}}{\underbrace{\left(\frac{1}{T}\sum_{t=1}^{T}\bm{S}_{t}\bm{\Gamma}\bm{A}^{-\top}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})+\frac{1}{T}\sum_{t=1}^{T}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{A}^{-1}\bm{\Gamma}\bm{S}_{t}^{\top}\right)}}.$
@@ A.2.1 Berry-Esseen bounds
The following theorem, due to Shao and Zhang (2022), gives a Berry-Esseen bound for non-linear statistics which would be crucial for our analysis of the non-asymptotic distribution of the TD estimation error.
@@
(Theorem 2.1 of Shao and Zhang (2022)) Let $\bm{X}_{1},...,\bm{X}_{n}$ be independent random variables, and let $\bm{T}(\bm{X}_{1},...,\bm{X}_{n})$ be a $d$-dimensional statistic that can be represented as
@@ A.2.2 Properties of multivariate Gaussian distributions
The following theorem by Devroye et al. (2018) gives an upper bound for the total variation (TV) distance between two Gaussian random variables with same means and different covariance matrices.
Theorem A.14 (Devroye et al. (2018)).
@@ A.3 Concentration inequalities
The following Azuma’s inequality for matrix martingales is shown in Tropp (2011).
@@
Specifically, for vector-valued martingales, the following proposition is shown by Kallenberg and Sztencel (1991).
@@
$W_{\max}\geq\sum_{i=1}^{t}\|\bm{x}_{i}-\bm{x}_{i-1}\|_{2}^{2}\quad\text{almost surely.}$.
@@
The following Freedman’s inequality on Hilbert spaces was shown by Peng et al. (2024) in their Theorem 2.1.
@@ B.1.1 High-probability convergence guarantee for the original TD estimation error
$16\eta_{0}\sqrt{\frac{2}{2\alpha-1}}\sqrt{\log\frac{9t^{\star}}{\delta}}\leq 1$ (71)
@@
$\mathcal{H}_{i}=\{\max_{1\leq j\leq i}\|\bm{\Delta}_{j}\|\leq R\},\quad\text{and}\quad\tilde{\bm{\zeta}}_{i}=\bm{\zeta}_{i}\mathds{1}(\mathcal{H}_{i-1}).$
@@
$\|\tilde{\bm{\zeta}}_{i}\|_{2}\leq 4R+2\|\bm{\theta}^{\star}\|_{2}+1.$
@@
Since $\bm{\theta}_{0}=0$, the norm of $\bm{\Delta}_{0}$ is equal to that of $\bm{\theta}^{\star}$, and therefore $\mathbb{P}(\mathcal{H}_{0})=1$ is trivial. When (72) holds for $t-1$, we consider the probability of $\mathcal{H}_{t}$. We firstly observe that
$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\tilde{\bm{\zeta}}_{i}\right\|_{2}^{2}$ $\leq\sum_{i=1}^{t}\|\bm{R}_{i}^{t}\|^{2}\|\tilde{\bm{\zeta}}_{i}\|_{2}^{2}$
@@
$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\tilde{\bm{\zeta}}_{i}\right\|_{2}$ $<2\eta_{0}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{\frac{2}{2\alpha-1}\log\frac{9t^{\star}}{\delta}}$
@@
$=\left\|\prod_{k=1}^{t}(\bm{I}-\eta_{k}\bm{A})\bm{\Delta}_{0}\right\|_{2}+\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\tilde{\bm{\zeta}}_{i}\right\|_{2}$
@@
$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\tilde{\bm{\zeta}}_{i}\right\|_{2}$ $\leq 2\eta_{0}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{2W_{t}\log\frac{3}{\delta_{t}}}$
@@
$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\tilde{\bm{\zeta}}_{i}\right\|_{2}\leq 2\eta_{0}\left(\frac{2\alpha}{\beta}\right)^{\frac{\alpha}{2(1-\alpha)}}\sqrt{\frac{2}{2\alpha-1}\log\frac{9T}{\delta}}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)t^{-\frac{\alpha}{2}}.$
@@
$\tilde{\bm{\Delta}}_{t}=\bm{\Delta}_{t}\mathds{1}(\mathcal{H}_{t}).$
@@
$\left\|\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\tilde{\bm{\Delta}}_{i-1}\right\|_{2}\leq 2\sqrt{2W_{\max}\log\frac{9}{\delta}},$
@@
$W_{\max}=\sup\frac{1}{T^{2}}\sum_{i=1}^{T}\|\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\tilde{\bm{\Delta}}_{i-1}\|_{2}^{2}$ $\lesssim\frac{1}{T^{2}}3\eta_{0}^{2}\left(\frac{2}{\beta}\right)^{\frac{1}{1-\alpha}}\cdot 4R^{2}\sum_{i=1}^{T}i^{-\alpha}$
@@
$\left\|\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\tilde{\bm{\Delta}}_{i-1}\right\|_{2}$ $\lesssim\eta_{0}\left(\frac{2}{\beta}\right)^{\frac{1}{2(1-\alpha)}}R\sqrt{\frac{1}{1-\alpha}\log\frac{9}{\delta}}$
@@
$\tilde{C}$ $=\eta_{0}^{2}\left(\frac{2}{\beta}\right)^{\frac{2+\alpha}{2(1-\alpha)}}(6\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{\frac{1}{(1-\alpha)(2\alpha-1)}}.$ (78)

### v4

@@ 3.4 Statistical inference
Combining the results from Sections 3.2 and 3.3, we describe a general methodology for constructing convex confidence sets for $\bm{\theta}^{\star}$ with guaranteed finite sample coverage. Based on the TD estimator $\overline{\bm{\theta}}_{T}$ in (10) and the variance estimator $\widehat{\bm{\Lambda}}_{T}$ (18), and given any tolerance level $\delta\in(0,1)$, it is enough to pick any convex set $\mathcal{C}_{\delta}$ that has probability at least $1-\delta$ under the Gaussian distribution $\mathcal{N}(\overline{\bm{\theta}}_{T},\frac{\widehat{\bm{\Lambda}}_{T}}{T})$ with random mean and covariance; that is, satisfying the condition
@@
Natural choices of $\mathcal{C}_{\delta}$ are (i) hyper-rectangles, leading to simultaneous confidence intervals of the form
$\mathcal{C}_{\delta}=\prod_{j=1}^{d}[(\overline{\bm{\theta}}_{T})_{j}-\widehat{R}/\sqrt{T},(\overline{\bm{\theta}}_{T})_{j}+\widehat{R}/\sqrt{T}],$ (23)
where $\widehat{R}>0$ is the (random) quantile of the $\ell_{\infty}$ norm of a centered Gaussian random vectors with covariance $(\widehat{\bm{\Lambda}}_{T}/{T})$, which can be calculated via simulations; and (ii) ellipsoids or balls, for example
@@
The next result, whose proof is in Appendix B.4, shows that the coverage of $\mathcal{C}_{\delta}$ converges to the nominal level of $1-\delta$ at a rate $O(T^{-1/3})$, as a function of $T$ alone.
@@
We highlight two important features of the above result. First, the coverage guarantee holds uniformly over all choices of the convex set $\mathcal{C}_{\delta}$ meeting condition (22), such as those in (23) and (24). Secondly, the size (i.e., volume) of the confidence sets vanishes at the standard (and optimal) parametric rate of $O(T^{-1/2})$, keeping all the other variables constant.
@@ 4 SIMULATIONS
We corroborate our theoretical findings with numerical experiments. Towards that goal, we consider a family of MDPs that are constructed to be difficult to distinguish from each other. The specification of this family of MDPs, along with detailed descriptions of our experiments, is provided in Appendix D. Here we summarize some of our results. We ran $10,000$ independent trials, each for $T=2^{20}$ iterations, and recorded data at iterations $2^{1},2^{2},2^{3},...,2^{19},2^{20}$. We report the rate of convergence of $\|\overline{\bm{\Delta}}_{T}\|_{2}^{2}$, the rate of converge of $\overline{\bm{\Delta}}_{T}$ to its asymptotic Gaussian approximation, and the precision of our estimator for the covariance matrix, under different choices of the parameter $\alpha=\frac{1}{2},\frac{2}{3},\frac{3}{4},1$. The results of these experiments confirm that the convergence speed is in accordance with our theoretical findings. In particular, for the choices of $\alpha=\frac{2}{3}$, the statistical inference procedure described in Section 3.4 achieves the target coverage rate when $T$ is sufficiently large.
Figure 2(a) illustrates the behavior of $\|\overline{\bm{\Delta}}_{T}\|_{2}^{2}$ under different choices of $\alpha$. As a reference, the black dotted line represents the $95\%$ quantile of $\{\bm{Z}_{T}\}$ where $\bm{Z}_{T}\sim\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}/T)$. Our theory predicts that when $T$ grows large, this should be the asymptotic line for the empirical $95\%$ quantiles when $\alpha\in(\frac{1}{2},1)$. This is indeed what is shown for $\alpha=\frac{2}{3}$ and $\frac{3}{4}$. When $\alpha=1$, however, the empirical quantiles do not appear to converge to the asymptotic quantiles. This fact is also consistent with our theory, as reflected in the $O(T^{\frac{\alpha}{2}-1})$ term in the bound of Theorem 3.1, which when $\alpha=1$, converges at the $O(T^{-\frac{1}{2}})$ rate, just like the leading term. As for the case of $\alpha=\frac{1}{2}$, we observe that when $T$ is small, $\|\overline{\bm{\Delta}}_{T}\|_{2}$ is relatively large and the empirical and asymptotic lines begin to overlap only when $T$ is very large (i.e., $T>10^{5}$). This phenomenon is also predicted by the $O(T^{-\frac{\alpha+1}{2}})$ term in the bound of Theorem 3.1, which vanishes at a slower rate when $\alpha=\frac{1}{2}$. Notice that for ease of computation, we use a different measure of the distributional difference than the convex distance, as specified in Eq. (124) in Appendix D.
Figure 2(b) depicts the speed of convergence of the empirical distribution of the rescaled TD estimation error $\sqrt{T}\overline{\bm{\Delta}}_{T}$ to its asymptotic distribution $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$. When $\alpha=1$, this distance does not converge to 0 as $T$ grows large, attesting to our claim that the $O(T^{\alpha-1})$ term in the distributional difference is inevitable. On the other hand, choosing $\alpha=\frac{1}{2}$, as done in Samsonov et al., 2024a , leads to a slower convergence than with $\alpha=\frac{2}{3}$ (our choice) and $\alpha=\frac{3}{4}$.
Figure 2: (a) Empirical $95\%$ quantiles of $\|\overline{\bm{\Delta}}_{T}\|_{2}$, and (b) distance between the empirical and asymptotic distributions of $\overline{\bm{\Delta}}_{T}$.
@@ 5 DISCUSSION
We develop high-probability convergence guarantees and Berry-Esseen bounds for TD learning with linear function approximation and $i.i.d.$ samples. We also propose an efficient way to construct a covariance estimator and derive its consistency rates. These results lead to a practicable methodology for constructing convex confidence sets for the target parameter, with its size vanishing at a parametric rate $O(T^{-1/2})$. An important direction for future research is to investigate whether the order $O(T^{-1/3})$ for the convergence speed of coverage rate in Theorem 3.4 is tight. Another interesting extension would be to derive specialized Berry-Esseen bounds for the sub-class of hyper-rectangles, where one would expect to find a better dependence on the dimension $d$ of the feature space, possibly of poly-logarithmic, as opposed to polynomial, order.
Our analysis and results can be generalized to cover other iterative algorithms, such as stochastic gradient descent and stochastic approximation methods. In Appendix E, we describe conditions under which our bounds and rates apply to the output of a general linear stochastic approximation algorithm for i.i.d. data, a setting that has received attention in the recent literature (see, e.g., Sheshukova et al.,, 2025; Samsonov et al., 2024a, ).
Acknowledgements
@@ A.1 Useful algebraic facts
For any complex number $z\in\mathbb{C}$, we use $\mathsf{Re}(z)$ to denote its real part. If a matrix $\bm{M}$ is positive (semi-)definite, i.e. $\lambda_{\min}(\bm{M})>(\geq)0$, we write $\bm{M}\succ(\succeq)\bm{0}$; when two matrices $\bm{X}$ and $\bm{Y}$ satisfy $\bm{X}-\bm{Y}\succ(\succeq)\bm{0}$, we write $\bm{X}\succ(\succeq)\bm{Y}$ or equivalently $\bm{Y}\prec(\preceq)\bm{X}$. For a vector $\bm{x}\in\mathbb{R}^{d}$, we use $\{x_{i}\}_{1\leq i\leq d}$ to denote its entries; for a matrix $\bm{X}\in\mathbb{R}^{m\times n}$, we use $\{X_{ij}\}_{1\leq i\leq m,1\leq j\leq n}$ to denote its entries. The vectorization of matrix $\bm{X}\in\mathbb{R}^{m\times n}$, denoted as $\textbf{vec}(\bm{X})$, is defined as a vector in $\mathbb{R}^{mn}$ with entries $\textbf{vec}(\bm{X})]_{(i-1)m+j}=\bm{X}_{ij}$ for every $1\leq i\leq m$ and $1\leq j\leq n$.
For a matrix sequence $\{\bm{F}_{t}\}_{t\in\mathbb{N}}$, a number sequence $\{f_{t}\}_{t\in\mathbb{N}}$ and a fixed matrix $\bm{X}$, if there exists a constant $\widetilde{C}$ independent of $t$ and $\bm{X}$ such that
@@
The following lemma, due to Chen et al., (2020)[Lemma 4.1], bounds the difference between the inverse matrices and would be useful for our analysis of the plug-in variance estimator.
@@
$0$ $\leq\mathbb{E}[f^{2}(s)]-2\mathbb{E}[f(s)\bm{\phi}^{\top}(s)\bm{y}]+\bm{y}^{\top}\mathbb{E}[\bm{\phi}(s)\bm{\phi}(s)^{\top}]\bm{y}$
@@
$\overset{(i)}{<}C+C\left(\frac{k+1}{k}\right)^{\nu-\alpha}\cdot\frac{\nu-\alpha}{k+1}-\beta C(k+1)^{-\alpha}\left(\frac{k+1}{k}\right)^{\nu-\alpha}+(k+1)^{-\alpha}$
$\overset{(ii)}{<}C-\frac{\beta}{2}C(k+1)^{-\alpha}\left(\frac{k+1}{k}\right)^{\nu-\alpha}+(k+1)^{-\alpha}$
$\overset{(iii)}{<}C-(k+1)^{-\alpha}\left(\frac{k+1}{k}\right)^{\nu-\alpha}+(k+1)^{-\alpha}<C.$
@@
where $Z_{1}=1^{\alpha}X_{1}-\frac{1}{\beta}=1-\frac{1}{\beta}$. The proof now boils down to showing that both $I_{1}$ and $I_{2}$ converges to 0 as $t\to\infty$. On one hand, $I_{1}$ is bounded by
@@
$\overset{(i)}{=}\frac{1}{\beta}\left(1+\sum_{i=t}^{\infty}\prod_{k=t+1}^{i+1}(1-\beta k^{-\alpha})-\sum_{j=t}^{\infty}\prod_{k=t+1}^{j+1}(1-\beta k^{-\alpha})\right)=\frac{1}{\beta}.$ (48)
@@
$-\underset{I_{23}}{\underbrace{\left(\frac{1}{T}\sum_{t=1}^{T}\bm{S}_{t}\bm{\Gamma}\bm{A}^{-\top}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})+\frac{1}{T}\sum_{t=1}^{T}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{A}^{-1}\bm{\Gamma}\bm{S}_{t}^{\top}\right)}}.$
@@ A.2.1 Berry-Esseen bounds
The following theorem, due to Shao and Zhang, (2022), gives a Berry-Esseen bound for non-linear statistics which would be crucial for our analysis of the non-asymptotic distribution of the TD estimation error.
@@
(Theorem 2.1 of Shao and Zhang, (2022)) Let $\bm{X}_{1},...,\bm{X}_{n}$ be independent random variables, and let $\bm{T}(\bm{X}_{1},...,\bm{X}_{n})$ be a $d$-dimensional statistic that can be represented as
@@ A.2.2 Properties of multivariate Gaussian distributions
The following theorem by Devroye et al., (2018) gives an upper bound for the total variation (TV) distance between two Gaussian random variables with same means and different covariance matrices.
Theorem A.14 (Devroye et al., (2018)).
@@ A.3 Concentration inequalities
The following Azuma’s inequality for matrix martingales is shown in Tropp, (2011).
@@
Specifically, for vector-valued martingales, the following proposition is shown by Kallenberg and Sztencel, (1991).
@@
$W_{\max}\geq\sum_{i=1}^{t}\|\bm{x}_{i}-\bm{x}_{i-1}\|_{2}^{2}\quad\text{almost surely.}$
@@
The following Freedman’s inequality on Hilbert spaces was shown by Peng et al., (2024) in their Theorem 2.1.
@@ B.1.1 High-probability convergence guarantee for the original TD estimation error
$16\eta_{0}\sqrt{\frac{2}{2\alpha-1}}\sqrt{\log\frac{9t^{\star}}{\delta}}\leq 1.$ (74)
@@
$\mathcal{H}_{i}=\{\max_{1\leq j\leq i}\|\bm{\Delta}_{j}\|\leq R\},\quad\text{and}\quad\widetilde{\bm{\zeta}}_{i}=\bm{\zeta}_{i}\mathds{1}(\mathcal{H}_{i-1}).$
@@
$\|\widetilde{\bm{\zeta}}_{i}\|_{2}\leq 4R+2\|\bm{\theta}^{\star}\|_{2}+1.$
@@
Since $\bm{\theta}_{0}=0$, the norm of $\bm{\Delta}_{0}$ is equal to that of $\bm{\theta}^{\star}$, and therefore $\mathbb{P}(\mathcal{H}_{0})=1$ is trivial. When (75) holds for t1t-1, we consider the probability of $\mathcal{H}_{t}$. We firstly observe that
$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\widetilde{\bm{\zeta}}_{i}\right\|_{2}^{2}$ $\leq\sum_{i=1}^{t}\|\bm{R}_{i}^{t}\|^{2}\|\widetilde{\bm{\zeta}}_{i}\|_{2}^{2}$
@@
$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\widetilde{\bm{\zeta}}_{i}\right\|_{2}$ $<2\eta_{0}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{\frac{2}{2\alpha-1}\log\frac{9t^{\star}}{\delta}}$
@@
$=\left\|\prod_{k=1}^{t}(\bm{I}-\eta_{k}\bm{A})\bm{\Delta}_{0}\right\|_{2}+\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\widetilde{\bm{\zeta}}_{i}\right\|_{2}$
@@
$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\widetilde{\bm{\zeta}}_{i}\right\|_{2}$ $\leq 2\eta_{0}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{2W_{t}\log\frac{3}{\delta_{t}}}$
@@
$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\widetilde{\bm{\zeta}}_{i}\right\|_{2}\leq 2\eta_{0}\left(\frac{2\alpha}{\beta}\right)^{\frac{\alpha}{2(1-\alpha)}}\sqrt{\frac{2}{2\alpha-1}\log\frac{9T}{\delta}}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)t^{-\frac{\alpha}{2}}.$
@@
$\widetilde{\bm{\Delta}}_{t}=\bm{\Delta}_{t}\mathds{1}(\mathcal{H}_{t}).$
@@
$\left\|\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\widetilde{\bm{\Delta}}_{i-1}\right\|_{2}\leq 2\sqrt{2W_{\max}\log\frac{9}{\delta}},$
@@
$W_{\max}=\sup\frac{1}{T^{2}}\sum_{i=1}^{T}\|\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\widetilde{\bm{\Delta}}_{i-1}\|_{2}^{2}$ $\lesssim\frac{1}{T^{2}}3\eta_{0}^{2}\left(\frac{2}{\beta}\right)^{\frac{1}{1-\alpha}}\cdot 4R^{2}\sum_{i=1}^{T}i^{-\alpha}$
@@
$\left\|\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\widetilde{\bm{\Delta}}_{i-1}\right\|_{2}$ $\lesssim\eta_{0}\left(\frac{2}{\beta}\right)^{\frac{1}{2(1-\alpha)}}R\sqrt{\frac{1}{1-\alpha}\log\frac{9}{\delta}}$
@@
In the mean time, since the difference between $\overline{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$ is bounded by
$\mathsf{Tr}(\overline{\bm{\Lambda}}_{T})-\mathsf{Tr}(\bm{\Lambda}^{\star})\lesssim T^{\alpha-1}\mathsf{Tr}(\bm{X}(\bm{\Lambda}^{\star}))\lesssim\frac{\mathsf{Tr}(\bm{\Lambda}^{\star})}{\beta}T^{\alpha-1},$
@@
$\widetilde{C}$ $=\eta_{0}^{2}\left(\frac{2}{\beta}\right)^{\frac{2+\alpha}{2(1-\alpha)}}(6\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{\frac{1}{(1-\alpha)(2\alpha-1)}}.$ (81)